OpenStudy (anonymous):

Is the equation y = x3 - 6 even, odd, or neither? A. even B. odd C. neither

OpenStudy (anonymous):

plug -x in the equation see what you get

OpenStudy (anonymous):

y=3-6-x or y=-3-x?

OpenStudy (anonymous):